OpenStudy (anonymous):

how do you answer this? tod wants to make a 12fl. oz of a 20% sugar solution by mixing together a 50% sugar solution and a 10% sugar solution. how much of each solution must he use?

OpenStudy (anonymous):

12 ounces at 20 percent means you want 20% of 12 = 2.4 ounces of sugar. if you put in x ounces of 50% solution it will have .5x sugar and you will then add 12 -x of 10% sugar giving .1(12-x) sugar. all together you will have \[.5x+.10(12-x)\] sugar and you want that to be 2.4 so solve \[.5x+.1(12-x)=2.4\] for x

OpenStudy (anonymous):

i would first multiply by ten to get rid of annoying decimals and write \[5x+(12-x)=24\] \[4x+12=24\] \[4x=12\] \[x=3\]

OpenStudy (anonymous):

Recall that x is the amount of the 50% solution and (12 - x), in this case 9, is the amount of 10% solution.

OpenStudy (anonymous):

where does the 12-x come from

OpenStudy (anonymous):

If you add x Oz of one fluid and want to fill a 12 Oz container, you need the difference between 12 and x, e.g. (12-x), Oz of the other fluid.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

ok i still dont get it

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

where the 12-x comes from i dont know if im doing right but i have an equation that looks like this .5x+.2x=3.6 and i dont know where to go from that

OpenStudy (anonymous):

i mean .5x+.1x=2.4

OpenStudy (anonymous):

was that provided to you or did you come up with that? because that would assume that you are adding equal amounts of each solution., which is not the case.